A Bibliography of IEEE Standards

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

10 September 2021 Version 1.17

Title word cross-reference

003.0 [IEE95a]. 003.0-1995 [IEE95a].

1003.1 [IEE88, IEE94a, IEE94b, IEE94c]. 1003.1-1988 [IEE88, IEE88]. 1003.1-1990 [IEE94a, IEE94b, IEE94c]. 1003.1-2001 [IEE01a, IEE01d, IEE01c, IEE01b]. 1003.1-2017 [IEE17a]. 1003.1/2003.1/INT [IEE94a, IEE94b, IEE94c]. 1003.10 [IEE95b]. 1003.10-1995 [IEE95b]. 1003.13 [IEE98a]. 1003.13-1998 [IEE98a]. 1003.1d [IEE99a]. 1003.1d-1999 [IEE99a]. 1003.1g [IEE00a]. 1003.1g-2000 [IEE00a]. 1003.1j [IEE00b]. 1003.1j-2000 [IEE00b]. 1003.2 [IEE92a, IEE92b]. 1003.2-1992 [IEE93, IEE92a, IEE92b]. 1003.2-1992/INT [IEE92a, IEE92b].

1003.23 [IEE98b]. 1003.23-1998 [IEE98b]. 1003.2a [IEE93]. 1003.2a-1992 [IEE93]. 1003.2d [IEE94d]. 1003.2d-1994 [IEE94d]. 1003.5 [IEE99b]. 1003.5-1999 [IEE99b]. 1003.5b [IEE99d]. 1003.5b-1999 [IEE99d]. 1003.9 [IEE92c]. 1003.9-1992 [IEE92c]. 1178 [IEE91]. 1178-1990 [IEE91]. 13210 [IEE99c]. 1387.2 [IEE95c]. 1387.2-1995 [IEE95c]. 1387.3 [IEE96a]. 1387.3-1996 [IEE96a]. 14519 [IEE99d]. 16509 [IEE99e]. 1788-2015 [IEE15]. 1788.1-2017 [IEE17b]. 1987 [IEE87]. 1988 [IEE88]. 1990 [IEE91]. 1992 [IEE92a, IEE92b, IEE92d, IEE92c, IEE94a, IEE94b, IEE94c]. 1992/INT [IEE92a, IEE92b]. 1994 [IEE94d]. 1995 [IEE95a, IEE95b, IEE95c]. 1996 [IEE96b, IEE96a, IEE96c]. 1997 [IEE00e]. 1998 [IEE98b, IEE98a]. 1999 [IEE99b, IEE99a, IEE00d, IEE00c].

2 [IEE93, IEE99h]. 2000 [IEE00b, IEE00f, IEE00a]. 2000.1 [IEE00c]. 2000.1-1998 [IEE99e]. 2000.1-1999 [IEE00c]. 2000.2 [IEE00d]. 2000.2-1999 [IEE00d]. 2003 [IEE00e]. 2003-1997 [IEE99c, IEE00e]. 2003.1 [IEE92d, IEE94a, IEE94b, IEE94c]. 2003.1-1992 [IEE92d, IEE94a, IEE94b, IEE94c]. 2003.2 [IEE96c]. 2003.2-1996 [IEE96c]. 2020 [IEE20].

6 [IEE01a, IEE01d, IEE01c]. 60559 [IEE20]. 60559-2020 [IEE20].

754 [Hou19, Hou21]. 754-1985 [IEE85]. 754-2008 [IEE08]. 754-2019 [IEE19].

854 [IEE87]. 854-1987 [ANS87, IEE87].

9945 [IEE93]. 9945-2 [IEE93].

Account [IEE96a]. Ada [IEE99b, IEE99d]. Additional [IEE99a]. Administration [IEE95c, IEE96a]. Advanced [IEE00b]. AEP [IEE98a]. Amendment [IEE94d, IEE98c, IEE99b, IEE99a, IEE99h, IEE00f, IEE99f]. American [AI83, ANS87]. ANSI [IEE93, IEE96b, IEE99d, AI83, IEE85]. ANSI/IEEE [AI83, IEE85]. API [IEE92c, IEE96b, IEE98c, IEE99b, IEE99d, IEE99a, IEE99f, IEE00f]. Application [IEE92c, IEE95b, IEE95d, IEE95e, IEE96b, IEE98a, IEE98c, IEE99b, IEE99d, IEE99a, IEE00f]. Arithmetic [IEE15, IEE17b, ANS87, IEE19, IEE85, IEE87, IEE08, IEE20]. Available [IEE99f].

Base [IEE01a]. Based [IEE95b]. Batch [IEE94d]. Binary [IEE85]. Binding [IEE92c, IEE99b, IEE99d]. Books [Hou19, Hou21].

C [IEE96b, IEE98c, IEE98d, IEE99a, IEE99f, IEE99g, IEE00b, IEE00f, IEE98c]. Checkpoint [IEE98d]. Checkpoint/Restart [IEE98d]. Computer [IEE88, AI83]. Conformance [IEE92d, IEE96c, IEE99c, IEE00e, IEE00f].

D10 [IEE95d]. D12 [IEE95e, IEE99h]. D16 [IEE98c]. D2 [IEE98d]. D5 [IEE99f]. D6 [IEE99g]. Dec [IEE98c]. December [IEE92a]. Definitions [IEE01a]. Description [IEE92c]. Developing [IEE98b]. Draft [IEE95d, IEE95e, IEE99f].

Edition [IEE88, IEE92a, IEE92b, IEE94a, IEE94b, IEE94c, IEE96b, IEE99b]. Environment [IEE95a, IEE95b, IEE95d, IEE95e, IEE98b, IEE98a, IEE94d]. Environments [IEE88]. Extension [IEE99f, IEE00f]. Extensions [IEE99a, IEE00b, IEE99d].

Feb [IEE95d, IEE95e]. Floating [ANS87, IEE19, IEE20, IEE85, IEE87, IEE08]. Floating-Point [ANS87, IEE19, IEE85, IEE87, IEE08, IEE20]. FORTRAN [IEE92c].

Group [IEE96a]. Guide [IEE95a, IEE98b]. Guidelines [IEE99c, IEE00e].

History [Hou19, Hou21].

Identical [IEE95a]. IEC [IEE20, IEE93, IEE95a, IEE96b, IEE99d, IEE99e, IEE99c]. IEEE [AI83, ANS87, IEE20, IEE87, IEE88, IEE92a, IEE92b, IEE92d, IEE92c, IEE93, IEE94a, IEE94b, IEE94c, IEE94d, IEE95a, IEE95b, IEE95c, IEE95d, IEE96b, IEE96a, IEE96c, IEE98b, IEE98a, IEE99b, IEE99d, IEE99e, IEE99c, IEE00d, IEE00c, IEE00e, IEE00f, IEE17b, AI83, Hou19, Hou21, IEE19]. IEEE/ANSI [IEE93, IEE96b, IEE99d].

Implementations [IEE99c, IEE00e]. Independent [ANS87, IEE87, IEE99b, IEE00a]. Information [IEE92d, IEE92c, IEE93, IEE94d, IEE95b, IEE95c, IEE95d, IEE96b, IEE96a, IEE96c, IEE98a, IEE99b, IEE99d, IEE99c, IEE99a, IEE00d, IEE00c, IEE00e, IEE00f, IEE01a, IEE01d, IEE01c, IEE01b]. Informative [IEE01b]. INT [IEE88, IEE92a, IEE92b, IEE94a, IEE94b, IEE94c]. INT/1992 [IEE88]. Interactive [IEE95e]. Interface [IEE88, IEE92c, IEE93, IEE94d, IEE95c, IEE96b, IEE96a, IEE98c, IEE98d, IEE99b, IEE99d, IEE99a, IEE00f, IEE01a, IEE01d, IEE01c, IEE01b]. Interfaces [IEE92c, IEE99b, IEE99d, IEE00a, IEE01d, IEE92d, IEE99b]. International [IEE20]. Interpretations [IEE88, IEE92a, IEE92b, IEE94a, IEE94b, IEE94c]. Interval [IEE15, IEE17b]. ISO [IEE20, IEE93, IEE95a, IEE96b, IEE99d, IEE99e, IEE99c]. ISO/IEC [IEE93, IEE95a, IEE96b, IEE99d, IEE99e, IEE99c]. ISO/IEC/IEEE [IEE20]. Issue [IEE01a, IEE01d, IEE01c].

July [IEE99f]. June [IEE99h].

Language [IEE91, IEE92c, IEE96b, IEE98c, IEE98d, IEE99b, IEE99d, IEE99a, IEE99g, IEE00b, IEE00f, IEE99f, AI83].

March [IEE92b, IEE94a]. Measuring [IEE92d, IEE96c, IEE99c, IEE00e, IEE00f]. Method [IEE99c, IEE00e]. Methods [IEE92d, IEE96c, IEE00f, IEE00d]. Multiprocessor [IEE95d].

National [ANS87]. Nov [IEE98d].

October [IEE94b, IEE94c]. One [Hou19, Hou21]. Open [IEE95a, IEE98b]. Operating [IEE88, IEE93, IEE94d, IEE95c, IEE96b, IEE96a, IEE99a, IEE01a, IEE01d, IEE01c, IEE01b]. Organization [IEE98b].

OSE [IEE95a, IEE98b].

P1003.14 [IEE95d]. P1003.18 [IEE95e]. P1003.1a [IEE98c]. P1003.1h [IEE99f]. P1003.1m [IEE98d]. P1003.1q [IEE99g]. P1003.2b [IEE99h]. P2003.1b [IEE00f]. P2003.1b-2000 [IEE00f]. Part [IEE92d, IEE92c, IEE93, IEE94d, IEE95c, IEE96b, IEE96a, IEE96c, IEE99b, IEE99a, IEE99f, IEE00f, IEE00a]. Pascal [AI83]. PII [IEE00a]. Point [ANS87, IEE19, IEE85, IEE87, IEE08, IEE20]. Portable [IEE88, IEE93, IEE94d, IEE95c, IEE96b, IEE96a, IEE99a, IEE01a, IEE01d, IEE01c, IEE01b]. POSIX [IEE01a, IEE01d, IEE01c, IEE01b, IEE92d, IEE92c, IEE93, IEE94d, IEE95a, IEE95b, IEE95d, IEE95e, IEE96b, IEE96a, IEE96c, IEE98a, IEE99b, IEE99d, IEE99c, IEE99a, IEE99f, IEE00e, IEE00f]. POSIX.1a [IEE17a]. Practice [IEE00d]. Profile [IEE95d, IEE98a, IEE95b, IEE95d, IEE95e]. Profiles [IEE98b]. Program [IEE92c, IEE96b, IEE98c, IEE99b, IEE99d, IEE99a, IEE00f]. Programming [IEE91, AI83]. Protocol [IEE99b, IEE00a].

R1994 [IEE87]. R1998 [IEE92c]. Radix [ANS87, IEE87]. Radix-Independent [ANS87, IEE87]. Rationale [IEE01b]. Realtime [IEE98a, IEE99d, IEE99a, IEE00b, IEE00f]. Recommended [IEE00d]. Reliable [IEE99f]. Requirements [IEE99c, IEE00e]. Restart [IEE98d].

Scheme [IEE91]. Sept [IEE99g]. Serviceable [IEE99f]. Services [IEE99f]. Shell [IEE93, IEE94d, IEE96c, IEE99h, IEE01c]. Simplified [IEE17b]. Specifications [IEE99c, IEE00e]. Standard [AI83, ANS87, Hou19, Hou21, IEE19, IEE20, IEE85, IEE87, IEE88, IEE92d, IEE92c, IEE94d, IEE95b, IEE95c, IEE95d, IEE96a, IEE96c, IEE98a, IEE99b, IEE99c, IEE00c, IEE00e, IEE00f, IEE01a, IEE01d, IEE01c, IEE01b, IEE15, IEE17b, IEE08, IEE91]. Standardized [IEE95d, IEE98a]. Standards [IEE92a, IEE92b, IEE94a, IEE94b, IEE94c, IEE99c, IEE00e]. Std [ANS87, IEE88, IEE91, IEE92a, IEE92b, IEE93, IEE94a, IEE94b, IEE94c, IEE96b, IEE99d, IEE99e, IEE99c, IEE01a, IEE01d, IEE01c, IEE01b, IEE17b, IEE17a]. Supercomputing [IEE95b]. Support [IEE98a]. System [IEE88, IEE92d, IEE92c, IEE93, IEE94d, IEE95a, IEE95c, IEE96b, IEE96a, IEE98b, IEE98c, IEE99b, IEE99d, IEE99a, IEE99f, IEE00f, IEE01a, IEE01d, IEE01c, IEE01b]. Systems [IEE95e, IEE99f].

Technology [IEE92d, IEE92c, IEE93, IEE94d, IEE95b, IEE95c, IEE95d, IEE96b, IEE96a, IEE96c, IEE98a, IEE99b, IEE99d, IEE99c, IEE99a, IEE00d, IEE00c, IEE00e, IEE00f, IEE01a, IEE01d, IEE01c, IEE01b]. Technology-Portable [IEE93]. Technology-Test [IEE00f]. Terminology [IEE99e, IEE00c]. Test [IEE92d, IEE96c, IEE99c, IEE00d, IEE00e, IEE00f]. TR [IEE95a]. Tracing [IEE99g].

Unapproved [IEE95d, IEE95e]. User [IEE96a, IEE98b]. Utilities [IEE94d, IEE99h, IEE01c, IEE93, IEE96c]. Utilities-Amendment [IEE94d].

x [IEE99a]. X3.97 [AI83]. X3.97-1983 [AI83]. xx [IEE00a].

Year [IEE99e, IEE00d, IEE00c].

References

ANSI:1983:AIX

[AI83] American National Standards Institute and Institute of Electrical and Electronics Engineers. ANSI/ IEEE 770 X3.97-1983: An American National Standard: IEEE Standard Pascal computer programming language. IEEE, New York, NY, USA, 1983. ISBN 0-471-88944-X. 128 pp. LCCN QA76.73.P2 A43 1983.

ANSI:1987:AIS

[ANS87] ANSI/IEEE. ANSI/IEEE Std 854-1987: An American National Standard: IEEE Standard for Radix-Independent Floating-Point Arithmetic. IEEE, New York, NY, USA, October 5, 1987. ISBN 0-7381-1167-

8. v + 14 pp. LCCN ???? US$44.00.

                                                URL http://ieeexplore.ieee.
                                                org/iel1/2502/1121/00027840.
                                                pdf;   http://ieeexplore.ieee.
                                                org/xpl/standardstoc.jsp?isnumber=
                                                1121&isYear=1987. Revised 1994.

INSPEC Accession Number: 3095617.

Hough:2019:ISO

[Hou19] David G. Hough. The IEEE Standard 754: One for the history books. Computer, 52(12): 109–112, December 2019. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                                URL http://grouper.ieee.org/
                                                groups/msc/ANSI_IEEE-Std-754-
                                                2019/background/ieee-computer.
                                                pdf;   https://www.computer.
                                                org/csdl/magazine/co/2019/12/
                                                08909942/1f8KFWxbTCU.

Hough:2021:ISO

[Hou21] David G. Hough. The IEEE Standard 754: One for the history books. ComputingEdge, ??(4):42– 46, August 2021. ISSN 2469-7087.

URL http://grouper.ieee.org/
groups/msc/ANSI_IEEE-Std-754-
2019/background/ieee-computer.
pdf;   https://www.computer.
org/csdl/magazine/co/2019/12/
08909942/1f8KFWxbTCU.

IEEE:1985:AIS

[IEE85] IEEE Task P754. ANSI/IEEE 754-1985, Standard for Binary Floating-Point Arithmetic. IEEE, New York, NY, USA, August 12, 1985. ISBN 1-55937-653-8. 20 pp. LCCN ???? US$35.00.

URL http://ieeexplore.ieee.
org/iel1/2355/1316/00030711.
pdf;   http://standards.ieee.
org/reading/ieee/std/busarch/
754-1985.pdf;   http://standards.
ieee.org/reading/ieee/std_public/
description/busarch/754-1985_
desc.html;   http://www.iec.
ch/cgi-bin/procgi.pl/www/iecwww.
p?wwwlang=E&wwwprog=cat-det.
p&wartnum=019113. Revised 1990.

A preliminary draft was published in the January 1980 issue of IEEE Computer, together with several companion articles [?, ?, ?, ?, ?, ?, ?]. The final version was republished in [?, ?]. See also [?]. Also standardized as IEC 60559 (1989-01) Binary floating-point arithmetic for microprocessor systems.

IEEE:1987:RIS

[IEE87] IEEE. 854-1987 (R1994) IEEE Standard for Radix-Independent Floating-Point Arithmetic. IEEE, New York, NY, USA, 1987. ISBN 1-55937-859-X. 16 pp. LCCN ???? US$44.00. URL http:

                                        //standards.ieee.org/reading/
                                        ieee/std/busarch/854-1987.pdf;
                                         http://standards.ieee.org/reading/
                                        ieee/std_public/description/
                                        busarch/854-1987_desc.html. Re-

vised 1994.

IEEE:1988:IEI

[IEE88] IEEE. 1003.1-1988 INT/1992 Edition, IEEE Standard Interpretations of IEEE Standard Portable Operating System Interface for Computer Environments (IEEE Std 1003.1-1988). IEEE, New York, NY, USA, 1988. ISBN 1-55937-216-8. 40 pp. LCCN ???? US$77.00.

IEEE:1991:ISI

[IEE91] IEEE. IEEE Std 1178-1990, IEEE Standard for the Scheme Programming Language. IEEE, New York, NY, USA, May 1991. ISBN 1-55937-125-0. 52 pp. LCCN ???? URL http:

                                        //standards.ieee.org/reading/
                                        ieee/std_public/description/
                                        busarch/1178-1990_desc.html.

IEEE:1992:IDE

[IEE92a] IEEE. 1003.2-1992/INT December 1994 Edition, IEEE Standards Interpretations for IEEE Std 1003.2-1992. IEEE, New York, NY, USA, 1992. ISBN 1-55937-476-4 (print), 0-7381-1375-1 (electronic). 64 pp. LCCN ???? US$71.00.

IEEE:1992:IME

[IEE92b] IEEE. 1003.2-1992/INT March 1994 Edition, IEEE Standards Interpretations for IEEE Std 1003.2-1992. IEEE, New York, NY, USA, 1992. ISBN 1-55937-411-X (print), 0-7381-1376-X (electronic). 40 pp. LCCN ???? US$92.00.

IEEE:1992:RIS

[IEE92c] IEEE. 1003.9-1992 (R1998) IEEE Standard for Information Technology — POSIX\begin{matrix}R \\ ◯ \end{matrix} FORTRAN 77 Language Interfaces — Part 1: Binding for System Application Program Interface (API) [Description]. IEEE, New York, NY, USA, 1992. ISBN 1-55937-230-3 (print), 0-7381-0635-6 (electronic). 186 pp. LCCN ???? US$98.00. URL http:

//standards.ieee.org/reading/
ieee/std_public/description/
posix/1003.9-1992_desc.html.

IEEE:1992:ISI

[IEE92d] IEEE. 2003.1-1992 IEEE Standard for Information Technology — Test Methods for Measuring Conformance to POSIX\begin{matrix}R \\ ◯ \end{matrix} — Part 1: System Interfaces. IEEE, New York, NY, USA, 1992. ISBN 1-55937-275-3 (print), 0-7381-0654-2 (electronic). 456 pp. LCCN ???? US$114.00. URL http:

//standards.ieee.org/reading/
ieee/std_public/description/
posix/2003.1-1992_desc.html.

IEEE:1993:III

[IEE93] IEEE. 9945-2: 1993 (ISO/IEC)

[IEEE/ANSI Std 1003.2-1992 and IEEE/ANSI 1003.2a-1992] Information Technology-Portable Operating System Interface (POSIX\begin{matrix}R \\ ◯ \end{matrix}) — Part 2: Shell and Utilities. IEEE, New York, NY, USA, 1993. ISBN 1-55937-406-3 (print), 0-7381-0657-7 (electronic). 1328 pp. LCCN ???? US$271.00.

IEEE:1994:IME

[IEE94a] IEEE. 1003.1/2003.1/INT March 1994 Edition, IEEE Standards Interpretations for IEEE Std 1003.1-1990 and IEEE Std 2003.1-1992. IEEE, New York, NY, USA, March 1994. ISBN 1-55937-410-1 (print), 0-7381-1373-5 (electronic). 48 pp. LCCN ???? US$61.00.

IEEE:1994:IOE

[IEE94b] IEEE. 1003.1/2003.1/INT, October 1994 Edition, IEEE Standards Interpretations for IEEE Std 1003.1-1990 and IEEE Std 2003.1-1992. IEEE, New York, NY, USA, October 1994. ISBN 1-55937-449-7 (print), 0-7381-1374-3 (electronic). 40 pp. LCCN ???? US$60.00.

IEEE:1994:IOEb

[IEE94c] IEEE. 1003.1/2003.1/INT, October 1994 Edition, IEEE Standards Interpretations for IEEE Std 1003.1-1990 and IEEE Std 2003.1-1992. IEEE, New York, NY, USA, October 1994. ISBN 1-55937-449-7 (print), 0-7381-1374-3 (electronic). 40 pp. LCCN ???? US$90.00.

IEEE:1994:ISI

[IEE94d] IEEE. 1003.2d-1994 IEEE Standard for Information Technology — Portable Operating System Interface (POSIX\begin{matrix}R \\ ◯ \end{matrix}) — Part 2: Shell and Utilities-Amendment 1: Batch Environment. IEEE, New York, NY, USA, 1994. ISBN 1-55937-512-4 (print), 0-7381-0633-X (electronic). 152 pp. LCCN ???? US$81.00.

IEEE:1995:IGP

[IEE95a] IEEE. 003.0-1995 IEEE Guide to the POSIX\begin{matrix}R \\ ◯ \end{matrix} Open System Environment (OSE) (Identical to ISO/IEC TR 14252). IEEE, New York, NY, USA, 1995. ISBN 1-55937-531-0. 288 pp. LCCN ???? US$98.00. URL http:

//standards.ieee.org/reading/
ieee/std_public/description/
posix/1003.0-1995_desc.html.

IEEE:1995:ISI

[IEE95b] IEEE. 1003.10-1995 IEEE Standard for Information Technology — POSIX\begin{matrix}R \\ ◯ \end{matrix} — Based Supercomputing Application Environment Profile. IEEE, New York, NY, USA, 1995. ISBN 1-55937-546-9 (print), 0-7381-0631-3 (electronic). 80 pp. LCCN ???? US$71.00. URL http:

//standards.ieee.org/reading/
ieee/std_public/description/
posix/1003.10-1995_desc.html.

IEEE:1995:ISIb

[IEE95c] IEEE. 1387.2-1995 IEEE Standard for Information Technology — Portable Operating Interface System — Part 2: Administration. IEEE, New York, NY, USA, 1995. ISBN 1-55937-537-X (print), 0-7381-0652-6 (electronic). 296 pp. LCCN ???? US$86.00. URL http:

//standards.ieee.org/reading/
ieee/std_public/description/
posix/1387.2-1995_desc.html.

IEEE:1995:UDP

[IEE95d] IEEE. Unapproved Draft P1003.14, D10 Feb 1995 IEEE Standard for Information Technology — POSIX\begin{matrix}R \\ ◯ \end{matrix} Standardized Profile — POSIX\begin{matrix}R \\ ◯ \end{matrix} Multiprocessor Application Environment Profile. IEEE, New York, NY, USA, February 1995. ISBN ???? ???? pp. LCCN ???? US$42.00.

IEEE:1995:UDPb

[IEE95e] IEEE. Unapproved Draft P1003.18, D12 Feb 1995 POSIX\begin{matrix}R \\ ◯ \end{matrix} Interactive Systems Application Environment Profile. IEEE, New York, NY, USA, February 1995. ISBN ???? ???? pp. LCCN ???? US$40.00.

IEEE:1996:ISI

[IEE96a] IEEE. 1387.3-1996 IEEE Standard for Information Technology — Portable Operating System Interface (POSIX\begin{matrix}R \\ ◯ \end{matrix}) System — Part 3: User and Group Account Administration. IEEE, New York, NY, USA, 1996. ISBN 1-55937-866-2 (print), 0-7381-0653-4 (electronic). 112 pp. LCCN ???? US$71.00.

IEEE:1996:III

[IEE96b] IEEE. 1996 (ISO/IEC) [IEEE/ ANSI Std 1003.1, 1996 Edition] Information Technology — Portable Operating System Interface (POSIX\begin{matrix}R \\ ◯ \end{matrix}) — Part 1: System Application: Program Interface (API) [C Language]. IEEE, New York, NY, USA, 1996. ISBN 1-55937-573-6. 784 pp. LCCN ???? US$143.00. URL http:

//standards.ieee.org/reading/
ieee/std_public/description/
posix/9945-1-1996_desc.html.

IEEE:1996:ISIb

[IEE96c] IEEE. 2003.2-1996 IEEE Standard for Information Technology — Test Methods for Measuring Conformance to POSIX\begin{matrix}R \\ ◯ \end{matrix} — Part 2: Shell and Utilities. IEEE, New York, NY, USA, 1996. ISBN 1-55937-882-4 (print), 0-7381-0655-0 (electronic). 1424 pp. LCCN ???? US$119.00.

IEEE:1998:ISI

[IEE98a] IEEE. 1003.13-1998 IEEE Standard for Information Technology — Standardized Application Environment Profile (AEP) — POSIX\begin{matrix}R \\ ◯ \end{matrix} Realtime Application Support. IEEE, New York, NY, USA, 1998. ISBN 0-7381-0178-8 (print), 0-7381-0632-1 (electronic). 186 pp. LCCN ???? US$77.00.

IEEE:1998:IGD

[IEE98b] IEEE. 1003.23-1998 IEEE Guide for Developing User Organization Open System Environment (OSE) Profiles. IEEE, New York, NY, USA, 1998. ISBN 0-7381-1541-X (print), 0-7381-1542-8 (electronic). 120 pp. LCCN ???? US$90.00.

IEEE:1998:PDD

[IEE98c] IEEE. P1003.1a, D16 Dec 98 System Application Program Interface (API) [C Language] — Amendment C. IEEE, New York, NY, USA, 1998. ???? pp. US$54.00.

IEEE:1998:PDN

[IEE98d] IEEE. P1003.1m, D2 Nov 1998 Checkpoint/Restart Interface [C Language]. IEEE, New York, NY, USA, November 1998. ???? pp. US$38.00.

IEEE:1999:ITP

[IEE99a] IEEE. 1003.1d-1999 Information Technology — Portable Operating System Interface (POSIX\begin{matrix}R \\ ◯ \end{matrix}) — Part 1: System Application Program Interface (API) — Amendment x: Additional Realtime Extensions [C Language]. IEEE, New York, NY, USA, 1999. ???? pp. US$57.00.

IEEE:1999:EIS

[IEE99b] IEEE. 1003.5-1999 Edition IEEE Standard for Information Technology — POSIX\begin{matrix}R \\ ◯ \end{matrix} Ada Language Interfaces — Part 1: Binding for System Application Program Interface (API) — Amendment 2: Protocol - Independent Interfaces. IEEE, New York, NY, USA, 1999. ISBN 0-7381-1539-8 (print), 0-7381-1540-1 (electronic). 890 pp. LCCN ???? US$145.00. This edition incorporates IEEE Std 1003.5-1992, IEEE Std 1003.5b-1996 and IEEE Std 1003.5c-1998.

IEEE:1999:IIIc

[IEE99c] IEEE. 13210:1999 (ISO/IEC) IEEE Std 2003-1997 IEEE Standard for Information Technology — Requirements and Guidelines for Test Method Specifications and Test Method Implementations for Measuring Conformance to POSIX\begin{matrix}R \\ ◯ \end{matrix} Standards. IEEE, New York, NY, USA, 1999. ISBN 0-7381-1797-8 (print), 0-7381-1798-6 (electronic). 96 pp. LCCN ???? US$86.00.

IEEE:1999:III

[IEE99d] IEEE. 14519:1999 1(ISO/IEC) IEEE/ANSI Std 1003.5b-1999 Information Technology — POSIX\begin{matrix}R \\ ◯ \end{matrix} Ada Language Interfaces — Binding for System Application Program Interface (API) — Realtime Extensions. IEEE, New York, NY, USA, 1999. ISBN 0-7381-1570-3. 548 pp. LCCN ???? US$95.00. URL http:

//standards.ieee.org/reading/
ieee/interp/1003.5.html.

IEEE:1999:IIIb

[IEE99e] IEEE. 16509:1999 (ISO/IEC) IEEE Std 2000.1-1998 Year 2000 Terminology. IEEE, New York, NY, USA, 1999. ISBN 0-7381-1803-6. 16 pp. LCCN ???? US$49.00.

URL http://grouper.ieee.org/
groups/2000/.

IEEE:1999:PDJ

[IEE99f] IEEE. P1003.1h D5, July 1999 Draft POSIX\begin{matrix}R \\ ◯ \end{matrix} Part 1: System API Extension — Services for Reliable, Available, and Serviceable Systems [C language] — amendment. IEEE, New York, NY, USA, July 1999. ???? pp. US$41.00.

IEEE:1999:PDS

[IEE99g] IEEE. P1003.1q, D6 Sept. 1999 Tracing [C Language]. IEEE, New York, NY, USA, September 1999. 124 pp. US$50.00.

IEEE:1999:PDJb

[IEE99h] IEEE. P1003.2b, D12 June 1999 Shell & Utilities, Amendment 2. IEEE, New York, NY, USA, June 1999. ???? pp. US$73.00.

IEEE:2000:PXP

[IEE00a] IEEE. 1003.1g-2000 Part xx: Protocol Independent Interfaces (PII). IEEE, New York, NY, USA, 2000. ISBN 0-7381-1888-5. ???? pp. LCCN ???? US$92.00.

IEEE:2000:ARE

[IEE00b] IEEE. 1003.1j-2000 Advanced Realtime Extensions [C Language]. IEEE, New York, NY, USA, 2000. ISBN 0-7381-1889-3. 100 pp. LCCN ???? US$53.00.

IEEE:2000:ISI

[IEE00c] IEEE. 2000.1-1999 IEEE Standard for Information Technology — Year 2000 Terminology. IEEE, New York, NY, USA, 2000. ISBN ???? 16 pp. LCCN ???? US$62.00.

IEEE:2000:IRP

[IEE00d] IEEE. 2000.2-1999 IEEE Recommended Practice for Information Technology — Year 2000 Test Methods. IEEE, New York, NY, USA, 2000. ISBN 0-7381-1779-X (print), 0-7381-1780-3 (electronic). 126 pp. LCCN ???? US$62.00.

URL http://grouper.ieee.org/
groups/2000/.

IEEE:2000:ISIb

[IEE00e] IEEE. 2003-1997 IEEE Standard for Information Technology — Requirements and Guidelines for Test Method Specifications and Test Method Implementations for Measuring Conformance to POSIX\begin{matrix}R \\ ◯ \end{matrix} Standards. IEEE, New York, NY, USA, 2000. ISBN 1-55937-895-6 (print), 0-7381-0240-7 (electronic). 96 pp. LCCN ???? US$71.00.

IEEE:2000:PIS

[IEE00f] IEEE. P2003.1b-2000 IEEE Standard for Information Technology-Test Methods for Measuring Conformance to POSIX\begin{matrix}R \\ ◯ \end{matrix} — Part 1:

System Application Program Interface (API) — Amendment 1: Realtime Extension [C Language]. IEEE, New York, NY, USA, 2000. ISBN 0-7381-2458-3. 394 pp. LCCN ???? US$77.00.

IEEE:2001:ISRa

[IEE01a] IEEE. IEEE Std 1003.1-2001 Standard for Information Technology — Portable Operating System Interface (POSIX) Base Definitions, Issue 6. IEEE, New York, NY, USA, 2001. ISBN 1-85912-247-7 (UK), 1-931624-07-0 (US), 0-7381-3047-8 (print), 0-7381-3010-9 (PDF), 0-7381-3129-6 (CD-ROM). xliv + 448 pp. LCCN ???? Revision of IEEE Std 1003.1-1996 and IEEE Std 1003.2-1992, Open Group Technical Standard Base Specifications, Issue 6.

IEEE:2001:ISRd

[IEE01b] IEEE. IEEE Std 1003.1-2001 Standard for Information Technology — Portable Operating System Interface (POSIX) Rationale (Informative). IEEE, New York, NY, USA, 2001. ISBN 1-85912-247-7 (UK), 1-931624-07-0 (US), 0-7381-3048-6 (print), 0-7381-3010-9 (PDF), 0-7381-3129-6 (CD-ROM). xxxiv + 310 pp. LCCN ???? Revision of IEEE Std 1003.1-1996 and IEEE Std 1003.2-1992) Open Group Technical Standard Base Specifications, Issue 6.

IEEE:2001:ISRc

[IEE01c] IEEE. IEEE Std 1003.1-2001 Standard for Information Technology — Portable Operating System Interface (POSIX) Shell and Utilities, Issue 6. IEEE, New York, NY, USA, 2001. ISBN 1-85912-247-7 (UK), 1-931624-07-0 (US), 0-7381-3050-8 (print), 0-7381-3010-9 (PDF), 0-7381-3129-6 (CD-ROM). xxxii + 1090 pp. LCCN ???? Revision of IEEE Std 1003.1-1996 and IEEE Std 1003.2-1992) Open Group Technical Standard Base Specifications, Issue 6.

IEEE:2001:ISRb

[IEE01d] IEEE. IEEE Std 1003.1-2001 Standard for Information Technology — Portable Operating System Interface (POSIX) System Interfaces, Issue 6. IEEE, New York, NY, USA, 2001. ISBN 1-85912-247-7 (UK), 1-931624-07-0 (US), 0-7381-3094-4 (print), 0-7381-3010-9 (PDF), 0-7381-3129-6 (CD-ROM). xxx + 1690 pp. LCCN ???? Revision of IEEE Std 1003.1-1996 and IEEE Std 1003.2-1992, Open Group Technical Standard Base Specifications, Issue 6.

P754:2008:ISF

[IEE08] IEEE Task P754. IEEE 754-2008, Standard for Floating-Point Arithmetic. IEEE, New York, NY, USA, August 29, 2008. ISBN 0-7381-5753-8 (paper), 0-7381-5752-X (electronic). 58 pp. LCCN ????

                                        URL http://en.wikipedia.org/
                                        wiki/IEEE_754-2008;   http://
                                        ieeexplore.ieee.org/servlet/
                                        opac?punumber=4610933.

IEEE:2015:ISI

[IEE15] IEEE. 1788-2015 — IEEE Standard for Interval Arithmetic. IEEE, New York, NY, USA, June 30, 2015. ISBN 0-7381-9721-1 (PDF), 0-7381-9720-3 (electronic). xiv + 79 pp. LCCN ???? URL http:/

/ieeexplore.ieee.org/servlet/
opac?punumber=7140719.  Ap-

proved 11 June 2015 by IEEE-SA Standards Board.

IEEE:2017:ISP

[IEE17a] IEEE. IEEE Std 1003.1-2017:

POSIX.1a. Web document, 2017.

URL https://pubs.opengroup.
org/onlinepubs/9699919799/. POSIX.1-

2017 is a revision of IEEE Std 1003.1-2008, and replaces previous versions of 2008, 2013, and 2016. Also issued as The Open Group Technical Standard Base Specifications, Issue 7.

IEEE:2017:ISI

[IEE17b] IEEE. IEEE Std 1788.1-2017 — IEEE Standard for Interval Arithmetic (Simplified). IEEE, New York, NY, USA, December 6, 2017. ISBN 1-5044-4603-8 (print), 1-5044-4602-X (PDF), 1-5044-4667-4. ix + 27 + 1 pp. LCCN ???? URL https:/

/standards.ieee.org/findstds/
standard/1788.1-2017.html.

IEEE-754:2019:ISF

[IEE19] IEEE-754. IEEE 754-2019, Standard for Floating-Point Arithmetic. IEEE, New York, NY, USA, June 13, 2019. ISBN 1-5044-5925-3 (print), 1-5044-5924-5 (e-PDF). 82 pp.

IEEE-754:2020:III

[IEE20] IEEE-754. 60559-2020 — ISO/IEC/IEEE International Standard — Floatingpoint arithmetic. IEEE, New York, NY, USA, 2020. ???? pp.

                                        URL https://standards.ieee.
                                        org/standard/60559-2020.html.